# Spectroscopy of two-dimensional interacting lattice electrons using symmetry-aware neural backflow transformations

### v1

@@ Acknowledgements.
We thank Dian Wu and Javier Robledo Moreno for engaging discussions. We especially thank Javier Robledo Moreno for sharing the data of Ref. [22]. We express our gratitude to Yusuke Nomura for providing valuable insights regarding quantum phase transitions and their connection to the structure of excitations. The open-source software NetKet version 3.9 was used to carry out the simulations [82, 83]. This work was supported by the Swiss National Science Foundation under Grant No. $200021\_200336$, and Microsoft Research.
Appendix A Numerical details and backflow architecture
@@
Given a set of occupation numbers, our symmetrized wavefunction ansatz is expressed as in Eq. (7), where we use a translation-invariant Jastrow factor defined in Eq. (3) with variational parameters $\theta_{J}=W_{d(ij)}$. The mean-field orbitals are captured by the matrix $M_{\mu,i}$ as defined in Eq. (1) of variational parameters of size $N_{f}\times N_{s}$. The backflow corrections are constructed as a translational equivariant convolutional neural network with parameters $\theta_{BF}$.
@@
The total set of variational parameters $\theta=(M_{\mu i},\theta_{BF},\theta_{J})$ are all optimized simultaneously using Variational Monte Carlo and Stochastic Reconfiguration [84]. To accelerate the optimization convergence of our Neural Slater-Backflow-Jastrow ansatz, we initialize the orbital parameters with converged parameters of a Hartree-Fock optimization.
@@
For smaller systems (i.e. $L=4$) we use shallow backflow CNN networks with one hidden layer and 32 features. For larger and more challenging systems such as $L=8,10$ we construct a residual CNN with 5 convolutional layers and 16 features per filter. Residual networks make the training more stable [85, 22, 51]. To access low-lying energy states, we use the latter residual CNN architecture regardless of the system size. Our deep CNNs use complex-valued weights for the feature maps. Each layer, except the final one, uses the complex Rectified Linear Unit (ReLU) activation function [86]. The output layer produces the backflow functions necessary for our computations.
@@
We optimize multiple CNN backflow transformation models in parallel, corresponding to the $N_{f}$ orbitals, ensuring efficient and scalable processing.
Appendix B Density-density Correlation Function
@@
To complement the results for $L=10$ shown in Fig. 4 (a), we also present the density-density correlation function, as defined in Eq. 11, for a system size of $L=8$ and $\bar{n}=0.44$ in Fig. 11. Similar to the $L=10$ case, we observe that the correlations display more pronounced oscillations as $V/t$ increases, indicating the increasing orderliness of the charge distribution during the phase transition. This behavior suggests a stronger tendency towards charge ordering with increasing interaction strength. In the charge-ordered (CO) phase, the amplitude of the oscillations gradually diminishes with distance. Conversely, in the metallic phase, the oscillations diminish more quickly with distance, highlighting a less ordered charge distribution. For weak couplings, correlations exhibit minimal oscillations and fade as the graph distance $R$ increases.
@@
Figure 11: The two-point density-density correlation functions as a function of graph distance $R$ for different values of $V/t$ for a system size of $L=8$ and $\bar{n}=0.44$.
Appendix C Excitation spectrum for closed-shell $L=10$ case
@@
In Fig. 12, we display the energy spectrum for $L=10$ and a closed-shell density $\bar{n}=0.41$, across different symmetry sectors for varying $V/t$. Similar to the $L=8$ case presented in the main text (Fig. 7), we observe that the ground state is situated within the $\Gamma$ sector.
@@
Figure 12: Lowest excitation energies in different $\mathbf{K}$ sectors for an $L=10$ closed-shell system with $\bar{n}=0.41$, comparing different VMC ansatz.
Appendix D Excitation Structure for Open-shell $L=4$ case
Moving away from closed shells induces changes in the structure of the excitation energies [87, 88, 89]. In Fig. 13, we benchmark our ansatz for the $L=4$ system with $\bar{n}=0.44$, which is not a closed-shell system. In the top panel, we compute the lowest and second-lowest energies using ED methods in each assigned $\mathbf{K}$ sector and compare the lowest energies in each sector to our VMC ansatzes. In the lower panel, we present the corresponding relative errors of the ansatzes. Our results show that the symmetric backflow improves the accuracy, demonstrating that our ansatz can also be effectively applied to non-closed-shell systems for computing excited states.
Figure 13: In the upper row, we compute the lowest and second-lowest energies for the designated $\mathbf{K}$ sector using ED and compare the lowest energies with our VMC ansatzes. The dashed lines in the lower row represent the relative errors of the ansatzes. These results correspond to an $L=4$ system for $\bar{n}=0.44$, which is not a closed-shell system.

### v2

@@ Acknowledgements.
We thank Dian Wu and Javier Robledo Moreno for engaging discussions. We especially thank Javier Robledo Moreno for sharing the data of Ref. Stokes_2020. We express our gratitude to Yusuke Nomura for providing valuable insights regarding quantum phase transitions and their connection to the structure of excitations. This work was supported by the Swiss National Science Foundation under Grant No. $200021\_200336$, and Microsoft Research.
Supplementary Note I Numerical details and backflow architecture
@@
Given a set of occupation numbers, our symmetrized wavefunction ansatz is expressed as in Eq. 12 of the main text, where we use a translation-invariant Jastrow factor defined in Eq. 8 of the main text, with variational parameters $\theta_{J}=W_{d(ij)}$. The mean-field orbitals are captured by the matrix $M_{\mu,i}$ as defined in Eq. 6 of the main text, of variational parameters of size $N_{f}\times N_{s}$. The backflow corrections are constructed as a translational equivariant convolutional neural network with parameters $\theta_{BF}$.
@@
The total set of variational parameters $\theta=(M_{\mu i},\theta_{BF},\theta_{J})$ are all optimized simultaneously using Variational Monte Carlo and Stochastic Reconfiguration Sorella_2001. To accelerate the optimization convergence of our Neural Slater-Backflow-Jastrow ansatz, we initialize the orbital parameters with converged parameters of a Hartree-Fock optimization.
@@
For smaller systems (i.e. $L=4$) we use shallow backflow CNN networks with one hidden layer and 32 features. For larger and more challenging systems such as $L=8,10$ we construct a residual CNN with 5 convolutional layers and 16 features per filter. Residual networks make the training more stable He_2015, Stokes_2020, Roth_2023. To access low-lying energy states, we use the latter residual CNN architecture regardless of the system size. Our deep CNNs use complex-valued weights for the feature maps. Each layer, except the final one, uses the complex Rectified Linear Unit (ReLU) activation function Trabelsi_2018. The output layer produces the backflow functions necessary for our computations.
@@
We optimize multiple CNN backflow transformation models in parallel, corresponding to the $N_{f}$ orbitals, ensuring efficient and scalable processing. For the optimization, we use a learning rate of 0.0075. The optimization is carried out using the SR preconditioner with a diagonal shift of 0.001, and we use 4096 Monte Carlo samples.
Supplementary Note II Density-density Correlation Function
@@
To complement the results for $L=10$ shown in Fig. 2 (a), we also present the density-density correlation function, as defined in Eq. 2 of the main text, for a system size of $L=8$ and $\bar{n}=0.44$ in Fig. S1. Similarly to the $L=10$ case, we observe that the correlations display more pronounced oscillations as $V/t$ increases, indicating the increasing orderliness of the charge distribution during the phase transition. This behavior suggests a stronger tendency towards charge ordering with increasing interaction strength. In the charge-ordered (CO) phase, the amplitude of the oscillations gradually diminishes with distance. Conversely, in the metallic phase, the oscillations diminish more quickly with distance, highlighting a less ordered charge distribution. For weak couplings, the correlations exhibit minimal oscillations and fade as the graph distance $R$ increases.
@@
Figure S1: Two-point correlations for $L=8$. The two-point density-density correlation functions as a function of graph distance $R$ for different values of interaction strength $V/t$ for a system size of $L=8$ and $\bar{n}=0.44$ with ground state momentum $\mathbf{K}=(\frac{3\pi}{4},0)$.
Supplementary Note III Excitation Spectrum for Closed-shell $L=10$ case
@@
In Fig. S2, we present the energy spectrum for $L=10$ and a closed-shell density $\bar{n}=0.41$, across different symmetry sectors for varying $V/t$. Similarly to the $L=8$ case discussed in the main text (Fig. 5), we observe that the ground state is located within the $\Gamma$ sector.
@@
Figure S2: Lowest excitation energies for $L=10$ closed-shell system. Lowest excitation energies in different $\mathbf{K}$ sectors for an $L=10$ closed-shell system with particle density $\bar{n}=0.41$, evaluated across various interaction strengths $V/t$. Results compare different variational ansatzes: symmetrized Neural Slater-Backflow-Jastrow ($\psi^{K}_{BF}$, blue) and symmetrized Slater-Jastrow without backflow ($\psi^{K}$, red). Error bars represent the corresponding standard deviations.
Supplementary Note IV Excitation Structure for Open-shell $L=4$ case
Moving away from closed shells induces changes in the structure of the excitation energies Dhar_2016, Brown_2017, Chiu_2019. In Fig. S3, we benchmark our ansatz for the $L=4$ system with $\bar{n}=0.44$, which is not a closed-shell system. In the upper panel, we compute the lowest and second-lowest energies using ED methods in each assigned $\mathbf{K}$ sector and compare the lowest energies in each sector to our VMC ansatzes. In the lower panel, we present the corresponding relative errors of the ansatzes. Our results show that the symmetric backflow improves the accuracy, demonstrating that our ansatz can also be effectively applied to non-closed-shell systems for computing excited states.
Figure S3: Energy comparison and relative errors for $L=4$ open-shell system. a) The lowest ($E_{0}^{K}$, teal horizontal lines) and second-lowest energies ($E_{1}^{K}$, yellow horizontal lines) for the designated $\mathbf{K}$ sector using exact diagonalization (ED), compared to the lowest energies in each sector obtained with our variational ansatzes: symmetrized Neural Slater-Backflow-Jastrow ($\psi^{K}_{BF}$, blue) and symmetrized Slater-Jastrow without backflow ($\psi^{K}$, red). b): The relative energy errors of the ansatzes with respect to the lowest energy in each sector are shown as dashed lines. These results correspond to a system with lattice size $L=4$ and density $\bar{n}=0.44$, which is not a closed-shell system. Error bars represent the corresponding standard deviations.
Supplementary Note V Effect of Network Size and Symmetry
In Fig. S4, we illustrate how the size of the neural network affects the accuracy of ground state energy accuracy for an $L=8$ system at a fixed $V/t$ interaction. The converged VMC energy is shown as a function of different layer sizes $\ell$ for various backflow neural network architectures, ranging from an equivariant CNN to a simple feed-forward neural network. The results indicate that increasing network size enhances energy accuracy. Dashed lines represent the no-backflow ansatzes: mean-field Slater determinant, Slater-Jastrow, and symmetrized Slater-Jastrow ($\psi^{K}$). In contrast, solid lines show the inclusion of backflow: a simple feedforward NN backflow and a symmetrized Neural Slater-Backflow-Jastrow with an equivariant CNN ($\psi^{K}_{BF}$). Additionally, we include data comparing a symmetry-aware convolutional neural network backflow ansatz that is not equivariant, where symmetry is incorporated by averaging over the symmetry group elements ($\psi^{K}_{BF}$-invariant), also shown with solid lines. The inset shows the results without the mean-field Slater ansatz, providing a zoomed-in view of the plot. The equivariant CNN-based neural network achieves lower energies than its invariant counterpart and other ansatzes. Although the invariant neural network backflow averages over group elements, it may not leverage the symmetry as effectively as the equivariant backflow. The equivariant neural network backflow integrates symmetry directly into each orbital’s backflow.
Figure S4: Impact of neural network size on energy accuracy. Effect of neural network size on ground state energy accuracy for a system of size $L=8$, with density $\bar{n}=0.44$ and ground state momentum $\mathbf{K}=(\frac{3\pi}{4},0)$ at fixed interaction strength $V/t=3.59$. Solid lines represent backflow-inclusive ansatzes, including equivariant and invariant neural networks. Dashed lines show no-backflow ansatzes. The inset provides a zoomed-in view of results without the mean-field Slater ansatz, showing the better performance of the equivariant CNN-based network in achieving lower energies. Error bars represent the corresponding standard deviations.